# Quantitative estimates for singularity for conjugate equations driven by linear fractional transformations

### v2

@@ 1. Introduction
The structure of this paper is as follows. In Section 2, we give some terminology and basic results of the conjugate equation on the unit interval. In Section 3, we deal with the singularity. Finally, we give an example in Section 4.
@@ 2.3. Linear fractional transforms
$h_{\sigma}(x)=h_{\sigma_{1}}\circ\cdots\circ h_{\sigma_{n}}(x),\ x\in[0,1].$
@@
Let $A_{i}:=\begin{pmatrix}a_{i}&b_{i}\\ c_{i}&d_{i}\end{pmatrix}$, $i\in\mathcal{I}_{N}$. We assume that $a_{i}d_{i}-b_{i}c_{i}=1$ for each $i$. Let $(i_{n})_{n}\in(\mathcal{I}_{N})^{\mathbb{N}}$. Let
@@ 3. Singularity
The main purpose of this section is to extend [27, Theorem 43] to the case that $f_{i}$ can be non-linear. In [27], the author dealt with many objects other than the conjugate functional equation, so the arguments were not specialized for the functional equation. Some parts dealing with the functional equation in [27] are somewhat sketchy, so we give more details here. We introduce some new notations and the formulations of our main results and some of the arguments in the proofs are different from [27, Theorem 43]. These changes will make arguments more transparent.
We assume that an LF-system $\{g_{i}\}_{i}$ is defined by $g_{i}(x):=\dfrac{a_{i}x+b_{i}}{c_{i}x+d_{i}},x\in[0,1]$, $i\in\mathcal{I}_{N}$, where $a_{i},b_{i},c_{i},d_{i},i\in\mathcal{I}_{N}$, are real numbers satisfying the conditions (1), (2) and (3) in Definition 2.7. In this section, we assume $d_{i}=1$ for each $i$.
@@
$\mathfrak{p}_{i}:=\left(G_{j}(\textup{Fix}(H_{i}))\right)_{j}\in\overline{\mathcal{P}_{N}},\ i\in\mathcal{I}_{N}.$
@@
If |$\left|\{\mathfrak{p}_{i}\}_{i\in\mathcal{I}_{N}}\right|\geq 2$, then, for every $\epsilon\in(0,1/2)$, there exists $\delta>0$ such that it holds that $\dim_{H}\mu_{\varphi}<1$ for every contractive system $(f_{i})_{i}$ on $[0,1]$ such that $\sum_{i}r_{i}<1+\delta$ and $\epsilon\leq r_{i}\leq 1-\epsilon$ for each $i$, where we let $r_{i}:=\textup{Lip}(f_{i})$.
The probability vectors $\mathfrak{p}_{i},i\in\mathcal{I}_{N}$, are determined only by $(\alpha,\beta,\{G_{i}\}_{i},\{H_{i}\}i)$ and date back to $\{g_{i}\}_{i}$. $\{f_{i}\}_{i}$ is not related to this. This contains the affine case that $f_{i}(x)=a_{i}x+(a_{0}+\cdots+a_{i-1})$ for each $i$. The reason why we have introduced $Y,G_{i}$ and $H_{i}$ will be clarified in the proof.
@@
If |$\left|\{\mathfrak{p}_{i}\}_{i\in\mathcal{I}_{N}}\right|=1$, then, for every $\epsilon\in(0,1/2)$, there exists $\delta>0$ such that it holds that $\dim_{H}\mu_{\varphi}<1$ for every contractive system $(f_{i})_{i}$ on $[0,1]$ such that $\sum_{i}r_{i}<1+\delta$, $\epsilon\leq r_{i}\leq 1-\epsilon$ for each $i$, and $\|{\bf r}-\mathfrak{p}_{0}\|\geq\epsilon$, where we let $r_{i}:=\textup{Lip}(f_{i})$ and ${\bf r}:=(r_{0},\cdots,r_{N-1})$.
If |$\left|\{\mathfrak{p}_{i}\}_{i\in\mathcal{I}_{N}}\right|=1$, then, $\mathfrak{p}_{0}=\cdots=\mathfrak{p}_{N-1}$. If $f_{i}(x)=a_{i}x+(a_{0}+\cdots+a_{i-1})$ for each $i$, and $\mathfrak{p}_{0}={\bf q}$, then, the solution $\varphi$ is smooth on $[0,1]$ and we also have an explicit expression for $\varphi$. See [27, Theorem 43 (ii)].
@@ 3.1. Proofs
Let $\epsilon\in(0,1/2)$. We assume that $\epsilon\leq\min_{i}r_{i}\leq\max_{i}r_{i}\leq 1-\epsilon$. Since $\{f_{i}\}_{i}$ is a compatible system, $\sum_{i}r_{i}\geq 1$ and $r_{i}>0$ for each $i$. Let $s\geq 1$ is a constant such that $\sum_{i\in\mathcal{I}_{N}}r_{i}^{s}=1$.
@@
it holds that $s<1+\dfrac{\epsilon_{2}}{\log 1/\epsilon}$ for every choice of $(r_{i})_{i}\in\mathcal{P}_{N}$ such that $\epsilon\leq r_{i}\leq 1-\epsilon$ for each $i$.
@@
This assertion is easily shown. If $\max_{k\in\mathcal{I}_{N}}|G_{k}(y)-q_{k}|>0$, then, $s_{N}\left((G_{k}(y))_{k}|{\bf q}\right)<0$. However, in our setting, we cannot exclude the possibility that $|G_{k}(y)-q_{k}|=0$ for every $k\in\mathcal{I}_{N}$. Assume that |$\left|\{\mathfrak{p}_{i}\}_{i}\right|\geq 2$. If $\max_{k\in\mathcal{I}_{N}}|G_{k}(y)-q_{k}|$ is close to 00, then, $\max_{k\in\mathcal{I}_{N}}|G_{k}(H_{i}(y))-q_{k}|$ is strictly positive. Hence, $W_{i}\left({\bf q},y\right)>0$.
@@
$p_{j}(0;z)=\left(G_{k}(H_{X_{j}(z)}\circ\cdots\circ H_{X_{1}(z)}(0))\right)_{k\in\mathcal{I}_{N}},$
@@
$p_{j+1}(0;z)=\left(G_{k}(H_{i_{1}}\circ H_{X_{j}(z)}\circ\cdots\circ H_{X_{1}(z)}(0))\right)_{k\in\mathcal{I}_{N}}.$
@@
If $(G_{j}(y))_{j}\notin B_{\mathscr{A}({\bf q})}$, then, $y\neq\mathfrak{p}_{\mathscr{A}({\bf q})}=\left(G_{j}(\textup{Fix}(H_{\mathscr{A}\left({\bf q}\right)}))\right)_{j}$ and hence, $y$ is not contained in an open interval $J_{\mathscr{A}({\bf q})}$ in $Y$ containing $\textup{Fix}(H_{\mathscr{A}\left({\bf q}\right)})$. Since $G_{j}$ and $H_{\mathscr{A}\left({\bf q}\right)}$ are continuous, $\max_{j}\left|G_{j}(H_{\mathscr{A}\left({\bf q}\right)}(y))-G_{j}(y)\right|$ is continuous with respect to $y$ on $Y$. Hence,
$\inf_{(G_{j}(y))_{j}\notin B_{\mathscr{A}({\bf q})}}\max_{j}\left|G_{j}(H_{\mathscr{A}\left({\bf q}\right)}(y))-G_{j}(y)\right|$
@@
Assume that $(G_{j}(y))_{j}\in B_{\mathscr{A}({\bf q})}$. Since ${\bf q}\notin B_{\mathscr{A}({\bf q})}$,
@@
Assume that $s\leq 1+\dfrac{\log\left(1+\epsilon/2\right)}{\log(1/\epsilon)}$. Since $\min_{i}r_{i}\geq\epsilon$, $\sum_{i}|r_{i}-r_{i}^{s}|\leq\epsilon^{1-s}-1\leq\epsilon/2$. Hence, $\left\|{\bf\widetilde{q}}-\mathfrak{p}_{0}\right\|\geq\epsilon/2$, where we let $\widetilde{q}_{i}:=r_{i}^{s}$ for each $i$ and ${\bf\widetilde{q}}:=(\widetilde{q}_{i})_{i}$. Since $\epsilon\leq 1/2$, $\widetilde{q}_{i}\geq\epsilon^{s}\geq\epsilon/2$ for each $i$. Hence, ${\bf\widetilde{q}}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$.
@@ 4. Example
We give an example that $\{g_{i}\}_{i}$ is an LF system and $\{f_{i}\}_{i}$ is non-affine and the solution $\varphi$ is singular. The result depends on the quantitative estimates in Section 3.
Let $N=2$. Then,
$G_{0}(y)=\frac{b_{1}(y+1)}{b_{1}y+1},\ H_{0}(y)=a_{0}y+c_{0},\ H_{1}(y)=\frac{(1-b_{1}+c_{1})y+c_{1}}{b_{1}y+1},$
@@
$g_{0}(x):=\frac{x}{-x+6},\ g_{1}(x):=\frac{3x+1}{5-x}.$
Then, $g_{0}$ is contractive, and $g_{1}$ is weakly contractive but not contractive. It holds that $a_{0}=1/6$, $c_{0}=-1/6$, $a_{1}=3/5$, $b_{1}=1/5$, and $c_{1}=-1/5$. Hence,
$\alpha=\min\{0,-1/5,-1\}=-1,\beta=\max\{0,-1/5,-1\}=0.$
@@
Let $\mathcal{P}[4/5]:=\left\{(p_{0},p_{1})\in\mathcal{P}_{2}:p_{0}\geq 4/5\right\}$. Let ${\bf q}=(q_{0},q_{1})\in\mathcal{P}[4/5]$. Then,
$|G_{0}(y)-q_{0}|=|G_{1}(y)-q_{1}|\geq\frac{3}{5},\ y\in[-1,0].$
@@
$s_{2}\left(p_{j}(0,z)|{\bf q}\right)\leq V\left({\bf q};\frac{6}{5}\right),\ j\geq 0,z\in(\mathcal{I}_{N})^{\mathbb{N}}.$ (4.1)
@@
$\limsup_{n\to\infty}\frac{\log R_{n}({\bf q};z)}{n}\leq-0.83,\ \textup{ $\nu_{0}$-a.s.$z$, \ ${\bf q}\in\mathcal{P}[4/5]$.}$ (4.2)
Let $\epsilon\in(0,1/2)$. Let $\{f_{0},f_{1}\}$ be a contractive system such that $\epsilon\leq r_{i}\leq 1-\epsilon$, where $r_{i}:=\textup{Lip}(f_{i})$, $i=0,1$. Let $s\geq 1$ be a constant such that $r_{0}^{s}+r_{1}^{s}=1$. Assume that $(r_{0}^{s},r_{1}^{s})\in\mathcal{P}[4/5]$. Then, by (4.2) and Lemma 3.5,
@@
We see that $\dfrac{0.83}{\log 10}\geq 0.36$ and $\dfrac{\log 2}{\log 2-\log 1.2}\leq 1.357$. Hence, (3.5) holds for $\epsilon=0.1$, $\epsilon_{2}=0.83$ and $\delta=0.2$. By Lemma 3.4, if $\{f_{0},f_{1}\}$ is a contractive system such that $0.1\leq r_{i}\leq 0.9$, $i=0,1$, $r_{0}+r_{1}\leq 1.2$, and $(r_{0}^{s},r_{1}^{s})\in\mathcal{P}[4/5]$, then,
@@
Now we can let $\{f_{0},f_{1}\}$ be a compatible system such that $0.85\leq r_{0}\leq 0.9$ and $0.1\leq r_{1}\leq 0.3$. Then, $r_{0}+r_{1}\leq 1.2$. By the compatibility condition, $r_{0}+r_{1}\geq 1$. The maps $f_{0}$ and $f_{1}$ may not be affine. Since $s\leq\frac{\log 2}{\log 2-\log 1.2}\leq 1.357$, $r_{0}^{s}\geq(0.85)^{1.357}\geq 0.8$. Hence, $(r_{0}^{s},r_{1}^{s})\in\mathcal{P}[4/5]$.
Remark 4.1.

### v3

@@ 1. Introduction
The structure of this paper is as follows. In Section 2, we give some terminology and basic results of the conjugate equation on the unit interval. In Section 3, we deal with the singularity. Finally, we give three numerical examples in Section 4.
@@ 2.3. Linear fractional transforms
$h_{\sigma}(x):=h_{\sigma_{1}}\circ\cdots\circ h_{\sigma_{n}}(x),\ x\in[0,1].$
@@
Let $A_{i}:=\begin{pmatrix}a_{i}&b_{i}\\ c_{i}&d_{i}\end{pmatrix}$, $i\in\mathcal{I}_{N}$. We assume that $a_{i}d_{i}-b_{i}c_{i}=1$ for each $i\in\mathcal{I}_{N}$. Let $(i_{n})_{n}\in(\mathcal{I}_{N})^{\mathbb{N}}$. Let
@@ 3. Singularity
The main purpose of this section is to extend [27, Theorem 43] to the case that $f_{i}$ can be non-linear. In [27], the author dealt with many objects other than the conjugate functional equation, so the arguments were not specialized for the functional equation. We introduce some new notations and the formulations of our main results and some of the arguments in the proofs are different from [27, Theorem 43]. These changes will make arguments more transparent.
We assume that an LF-system $\{g_{i}\}_{i}$ is defined by $g_{i}(x):=\dfrac{a_{i}x+b_{i}}{c_{i}x+d_{i}},x\in[0,1]$, $i\in\mathcal{I}_{N}$, where $a_{i},b_{i},c_{i},d_{i},i\in\mathcal{I}_{N}$, are real numbers satisfying the conditions (1), (2) and (3) in Definition 2.7. In this section, we assume $d_{i}=1$ for each $i\in\mathcal{I}_{N}$.
@@
${\bf G}(y):=(G_{0}(y),\dots,G_{N-1}(y))\in\overline{\mathcal{P}_{N}},\ y\in Y.$
Let
$\mathfrak{p}_{i}:={\bf G}\left(\textup{Fix}(H_{i})\right)\in\overline{\mathcal{P}_{N}},\ i\in\mathcal{I}_{N}.$
@@
If |$\left|\{\mathfrak{p}_{i}\}_{i\in\mathcal{I}_{N}}\right|\geq 2$, then, for every $\epsilon\in(0,1/2)$, there exists $\delta>0$ such that it holds that $\dim_{H}\mu_{\varphi}<1$ for every contractive system $(f_{i})_{i}$ on $[0,1]$ such that $\sum_{i\in\mathcal{I}_{N}}r_{i}<1+\delta$ and $\epsilon\leq r_{i}\leq 1-\epsilon$ for each $i\in\mathcal{I}_{N}$, where we let $r_{i}:=\textup{Lip}(f_{i})$.
The probability vectors $\mathfrak{p}_{i},i\in\mathcal{I}_{N}$, are determined only by $(\alpha,\beta,\{G_{i}\}_{i},\{H_{i}\}i)$ and date back to $\{g_{i}\}_{i}$. $\{f_{i}\}_{i}$ is not related to this. This contains the affine case that $f_{i}(x)=a_{i}x+(a_{0}+\cdots+a_{i-1})$ for each $i\in\mathcal{I}_{N}$. The reason why we have introduced $Y,G_{i}$ and $H_{i}$ will be clarified in the proof.
@@
If |$\left|\{\mathfrak{p}_{i}\}_{i\in\mathcal{I}_{N}}\right|=1$, then, for every $\epsilon\in(0,1/2)$, there exists $\delta>0$ such that it holds that $\dim_{H}\mu_{\varphi}<1$ for every contractive system $(f_{i})_{i}$ on $[0,1]$ such that $\sum_{i\in\mathcal{I}_{N}}r_{i}<1+\delta$, $\epsilon\leq r_{i}\leq 1-\epsilon$ for each $i\in\mathcal{I}_{N}$, and $\|{\bf r}-\mathfrak{p}_{0}\|\geq\epsilon$, where we let $r_{i}:=\textup{Lip}(f_{i})$ and ${\bf r}:=(r_{0},\cdots,r_{N-1})$.
If |$\left|\{\mathfrak{p}_{i}\}_{i\in\mathcal{I}_{N}}\right|=1$, then, $\mathfrak{p}_{0}=\cdots=\mathfrak{p}_{N-1}$. If $f_{i}(x)=a_{i}x+(a_{0}+\cdots+a_{i-1})$ for each $i\in\mathcal{I}_{N}$, and $\mathfrak{p}_{0}={\bf q}$, then, the solution $\varphi$ is smooth on $[0,1]$ and we also have an explicit expression for $\varphi$. See [27, Theorem 43 (ii)].
@@ 3.1. Proofs
Let $\epsilon\in(0,1/2)$. We assume that $\epsilon\leq\min_{i\in\mathcal{I}_{N}}r_{i}\leq\max_{i\in\mathcal{I}_{N}}r_{i}\leq 1-\epsilon$. Since $\{f_{i}\}_{i}$ is a compatible system, $\sum_{i\in\mathcal{I}_{N}}r_{i}\geq 1$ and $r_{i}>0$ for each $i\in\mathcal{I}_{N}$. Let $s\geq 1$ is a constant such that $\sum_{i\in\mathcal{I}_{N}}r_{i}^{s}=1$.
@@
it holds that $s<1+\dfrac{\epsilon_{2}}{\log 1/\epsilon}$ for every choice of $(r_{i})_{i\in\mathcal{I}_{N}}\in\mathcal{P}_{N}$ such that $\epsilon\leq r_{i}\leq 1-\epsilon$ for each $i\in\mathcal{I}_{N}$.
@@
This assertion is easily shown. If $\max_{k\in\mathcal{I}_{N}}|G_{k}(y)-q_{k}|>0$, then, $s_{N}\left({\bf G}(y)|{\bf q}\right)<0$. However, in our setting, we cannot exclude the possibility that $|G_{k}(y)-q_{k}|=0$ for every $k\in\mathcal{I}_{N}$. Assume that |$\left|\{\mathfrak{p}_{i}\}_{i}\right|\geq 2$. If $\max_{k\in\mathcal{I}_{N}}|G_{k}(y)-q_{k}|$ is close to 00, then, $\max_{k\in\mathcal{I}_{N}}|G_{k}(H_{i}(y))-q_{k}|$ is strictly positive. Hence, $W_{i}\left({\bf q},y\right)>0$.
@@
$p_{j}(0;z)={\bf G}\left(H_{X_{j}(z)}\circ\cdots\circ H_{X_{1}(z)}(0)\right),$
@@
$p_{j+1}(0;z)={\bf G}\left(H_{i_{1}}\circ H_{X_{j}(z)}\circ\cdots\circ H_{X_{1}(z)}(0)\right).$
@@
If ${\bf G}(y)\notin B_{\mathscr{A}({\bf q})}$, then, $y\neq\mathfrak{p}_{\mathscr{A}({\bf q})}={\bf G}\left(\textup{Fix}(H_{\mathscr{A}\left({\bf q}\right)})\right)$ and hence, $y$ is not contained in an open interval $J_{\mathscr{A}({\bf q})}$ in $Y$ containing $\textup{Fix}(H_{\mathscr{A}\left({\bf q}\right)})$. Since $G_{j}$ and $H_{\mathscr{A}\left({\bf q}\right)}$ are continuous, $\max_{j}\left|G_{j}(H_{\mathscr{A}\left({\bf q}\right)}(y))-G_{j}(y)\right|$ is continuous with respect to $y$ on $Y$. Hence,
$\inf_{{\bf G}(y)\notin B_{\mathscr{A}({\bf q})}}\max_{j}\left|G_{j}(H_{\mathscr{A}\left({\bf q}\right)}(y))-G_{j}(y)\right|$
@@
Assume that ${\bf G}(y)\in B_{\mathscr{A}({\bf q})}$. Since ${\bf q}\notin B_{\mathscr{A}({\bf q})}$,
@@
Assume that $s\leq 1+\dfrac{\log\left(1+\epsilon/2\right)}{\log(1/\epsilon)}$. Since $\min_{i}r_{i}\geq\epsilon$, $\sum_{i}|r_{i}-r_{i}^{s}|\leq\epsilon^{1-s}-1\leq\epsilon/2$. Hence, $\left\|{\bf\widetilde{q}}-\mathfrak{p}_{0}\right\|\geq\epsilon/2$, where we let $\widetilde{q}_{i}:=r_{i}^{s}$ for each $i\in\mathcal{I}_{N}$ and ${\bf\widetilde{q}}:=\left(\widetilde{q}_{i}\right)_{i\in\mathcal{I}_{N}}$. Since $\epsilon\leq 1/2$, $\widetilde{q}_{i}\geq\epsilon^{s}\geq\epsilon/2$ for each $i\in\mathcal{I}_{N}$. Hence, ${\bf\widetilde{q}}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$.
@@ 4. Example
We give three examples that $\{g_{i}\}_{i}$ is an LF system and $\{f_{i}\}_{i}$ is non-affine and the solution $\varphi$ is singular. The result depends on the quantitative estimates in Section 3.
We first deal with the case that $N=2$. Then,
$G_{0}(y)=\frac{b_{1}(y+1)}{b_{1}y+1},\ H_{0}(y)=a_{0}y+c_{0},\ H_{1}(y)=\frac{(1-b_{1}+c_{1})y+c_{1}}{b_{1}y+1}.$
Example 4.1.
@@
$g_{0}(x):=\frac{x}{-x+6},\textup{ and }g_{1}(x):=\frac{3x+1}{5-x}.$
Then, $g_{0}$ is contractive, and $g_{1}$ is weakly contractive but not contractive. We see that
$\alpha=\min\left\{0,-\frac{1}{5},-1\right\}=-1,\beta=\left\{0,-\frac{1}{5},-1\right\}=0.$
We see that
$G_{0}(y)=\frac{y+1}{y+5},\textup{ and }G_{1}(y)=1-G_{0}(y)=\frac{4}{y+5}.$
@@
It also holds that
$H_{0}(y)=\frac{y-1}{6}\textup{ and }H_{1}(y)=\frac{3y-1}{y+5}.$
Hence, $\textup{Fix}(H_{0})=\{-1/5\}$ and $\textup{Fix}(H_{1})=\{-1\}$. Hence,
${\bf G}(\textup{Fix}(H_{0}))=\left(\frac{1}{6},\frac{5}{6}\right)\textup{ and }{\bf G}(\textup{Fix}(H_{1}))=(0,1).$
Let $\widetilde{\mathcal{P}_{2}}:=\left\{(p_{0},p_{1})\in\mathcal{P}_{2}:p_{0}\geq 4/5\right\}$. Then,
$|G_{0}(y)-q_{0}|=|G_{1}(y)-q_{1}|\geq\frac{3}{5},\ y\in[-1,0],\ {\bf q}=(q_{0},q_{1})\in\widetilde{\mathcal{P}_{2}}.$
@@
$s_{2}\left({\bf G}(y)|{\bf q}\right)\leq V\left({\bf q};\frac{6}{5}\right),\ y\in[-1,0],\ {\bf q}\in\widetilde{\mathcal{P}_{2}}.$ (4.1)
@@
$\limsup_{n\to\infty}\frac{\log R_{n}({\bf q};z)}{n}\leq-0.83,\ \textup{ $\nu_{0}$-a.s.$z$, \ ${\bf q}\in\widetilde{\mathcal{P}_{2}}$.}$ (4.2)
Let $\epsilon\in(0,1/4)$. Let $\{f_{0},f_{1}\}$ be a contractive system such that $\epsilon\leq r_{i}\leq 1-\epsilon$, where $r_{i}:=\textup{Lip}(f_{i})$, $i=0,1$. Let $s\geq 1$ be a constant such that $r_{0}^{s}+r_{1}^{s}=1$. Assume that $(r_{0}^{s},r_{1}^{s})\in\widetilde{\mathcal{P}_{2}}$. Then, by (4.2) and Lemma 3.5,
@@
We see that $\dfrac{0.83}{\log 10}\geq 0.36$ and $\dfrac{\log 2}{\log 2-\log 1.2}\leq 1.357$. Hence, (3.5) holds for $N=2$, $\epsilon=0.1$, $\epsilon_{2}=0.83$ and $\delta=0.2$. By Lemma 3.4, if $\{f_{0},f_{1}\}$ is a contractive system such that $0.1\leq r_{i}\leq 0.9$, $i=0,1$, $r_{0}+r_{1}\leq 1.2$, and $(r_{0}^{s},r_{1}^{s})\in\widetilde{\mathcal{P}_{2}}$, then,
@@
Now we can let $\{f_{0},f_{1}\}$ be a compatible system such that $0.85\leq r_{0}\leq 0.9$, $0.1\leq r_{1}\leq 0.3$ and $r_{0}+r_{1}\geq 1$. Then, $r_{0}+r_{1}\leq 1.2$. The maps $f_{0}$ and $f_{1}$ can be non-affine. Since $s\leq\frac{\log 2}{\log 2-\log 1.2}\leq 1.357$, $r_{0}^{s}\geq(0.85)^{1.357}\geq 0.8$. Hence, $(r_{0}^{s},r_{1}^{s})\in\widetilde{\mathcal{P}_{2}}$.